# Fractality of Wireless Mesh Networks: Dimensional Effects on Network Performance

### v1

@@ Abstract
Wireless mesh networks (WMNs) depend on the spatial distribution of nodes, which directly influences connectivity, routing efficiency, and overall network performance. Conventional models typically assume uniform or random node placement, which inadequately represent the complex, hierarchical spatial patterns observed in practical deployments. In this study, we present a novel algorithm that constructs WMN topologies with tunable fractal dimensions, allowing precise control over spatial self-similarity. By systematically varying the fractal dimension, the algorithm generates network layouts spanning a continuum of spatial complexities, ranging from sparse fragmented clusters to dense, cohesive structures. Through NS-3 simulations, Key performance metrics including throughput, latency, jitter, and packet delivery ratio were evaluated across a range of fractal dimensions. Comparative evaluations against classical random, small-world, and scale-free network models reveal that high-dimensional fractal topologies achieve enhanced resilience and throughput under equivalent conditions. These findings demonstrate the potential of fractal geometry as a design paradigm for scalable and efficient WMN architectures.
@@ Results
Figure 1 illustrates that at a low fractal dimension ($D=1$), the network fragments into sparsely connected components, forming isolated clusters with weak interconnections. This fragmentation results from the inherent spatial sparsity caused by the recursive process of low-dimensional generation. As the fractal dimension increases, the network becomes increasingly integrated and spatially uniform. At $D=2$, the topology resembles a grid-like tiling, while at $D=3$, more cross-linked central paths emerge, forming distinctive cruciform structures. Interestingly, from $D\in[4,8]$, the generated networks converge toward a stable topological regime characterized by dense, isotropic interconnectivity and robust multi-hop pathways. This suggests a structural saturation effect, where further increases in fractal dimension yield diminishing changes in spatial arrangement. When $D=7$ and continuing through $D=9$, ring-shaped or circular substructures begin to appear within the network topology. These formations emergence of locally clustered subnetworks embedded within the overall topology, rather than a uniformly dispersed node distribution. This shift suggests that higher fractal dimensions promote localized aggregation, likely because of recursive spatial overlap during the point-generation process.
@@
Figure 3: Global efficiency and algebraic connectivity of networks with fractal dimensions varying from $D=1$ to $D=10$.
@@
Figure 5: Comparative performance analysis of fractal and baseline networks. Quality-of-service (QoS) metrics—including throughput, packet delivery ratio (PDR), end-to-end delay, and jitter are evaluated across fractal, scale-free (BA), small-world (WS), and random (ER) topologies under the same simulation conditions.
To reveal the structural origins of these performance trends, topological indicators including global efficiency and algebraic connectivity were further examined (Figure 3).Global efficiency, defined as the average of the inverse shortest path lengths between all node pairs, shows a non-monotonic pattern. it initially declines from $D=1$ to $D\approx 3$, then stabilizes for $D\geq 6$. Algebraic connectivity, representing spectral robustness and computed from the Laplacian matrix weighted by link capacity, grows steadily as $D$ increases, eventually plateauing beyond $D\geq 6$. These structural transitions shows the observed inflection in performance, suggesting that improved connectivity and reduced path fragmentation at higher fractal dimensions explain the improved communication.
@@
To further contextualize the advantages of high-dimensional fractal topologies, networks were compared with several classical baselines - including random geometric graphs, scale free and small-world configurations under equivalent traffic conditions (Figure 5). For this comparison, three representative fractal networks with dimensions $D=2$, $D=6.5$, and $D=7.5$ were selected. When $D=2$, the quasi-regular structure resembles grid-like arrangement. $D=6.5$ and $D=7.5$ correspond to configurations that achieved the highest performance on multiple metrics. The results confirm that fractal-based networks, particularly those at higher $D$ consistently perform better than baseline networks. These findings demonstrate that a higher fractal dimension yields a dual benefit: improved structural cohesion and enhanced performance.
@@ Discussion
This section presents an algorithm that constructs WMN topologies by explicitly setting the fractal dimension $D$ as an input parameter to control the spatial distribution of the nodes. The nodes are positioned according to the specified fractal dimension, and the edges of the network are formed by connecting pairs of nodes within a fixed communication radius $r$ The resulting fractal-based topologies are then analyzed using NS-3 simulations, where key performance indicators such as throughput, end-to-end delay, jitter, and PDR are measured under uniform traffic scenarios. To implement the proposed algorithm, the following steps are performed:
@@
$a_{ij}=\begin{cases}1,&\text{if node }i\text{ is directly connected to node }j,\\ 0,&\text{otherwise.}\end{cases}$ (1)
@@
1:Input: Target number of nodes $N$, fractal dimension $D$
2:Output: First $N$ node coordinates $(x_{i},y_{i})$
[…]
@@
[…]
@@
To establish a rigorous baseline for evaluating the structural and communication advantages of fractal-based networks, three classical generative network models were constructed: Erdős–Rényi (ER), Watts–Strogatz (WS), and Barabási–Albert (BA), all generated under consistent conditions of network size and mean degree. Specifically, ER networks were generated using the $G(n,p)$ model with $n=85$ nodes and edge probability $p=\frac{k}{n-1}$, targeting a mean degree $k=4.5$. WS networks were initialized as regular ring lattices with each node connected to its $k=4$ nearest neighbors and then rewired with a probability of $p=0.1$ to introduce small-world properties. BA networks were created using a preferential attachment mechanism, starting with a small fully connected seed and sequentially adding nodes, each establishing $m=2$ links based on the degree of the existing node. To ensure a valid comparison with the fractal networks, all generated networks were required to be fully connected, with repeated sampling applied as needed. Random seeds were fixed to maintain reproducibility and consistent layout configurations. These baseline models represent, respectively, random (ER), small-world (WS), and scale-free (BA) regimes commonly studied in network topology literature. They serve as essential structural archetypes for evaluating the topological distinctiveness and impact on performance of high-dimensional fractal geometries. The corresponding topologies are illustrated in Figure 7, spanning a spectrum from homogeneous randomness and modular shortcuts to centralized heterogeneity. This diversity provides a set of contrastive regimes for benchmarking the behavior of fractal-based network designs.
Figure 7: Visualization of classical network models (ER, WS, BA) used as baseline topologies for comparison. All configurations have identical node counts and mean degrees to ensure fair evaluation. Node colors represent average link capacity, highlighting local connectivity patterns.
@@
The simulation environment consisted of 85 static nodes uniformly deployed across a $500\,\mathrm{m}\times 500\,\mathrm{m}$ area. The node positions were generated using the proposed algorithm (Algorithm 1), with the fractal dimension $D\in[1,10]$. The resulting self-similar and spatial distributions were visually verified.
@@ Acknowledgements
We would like to express our sincerest gratitude to the Al-Farabi Kazakh National University for supporting this work by providing computing resources (Department of Physics and Technology). This research was funded by the Committee of Science of the Ministry of Education and Science of the Republic of Kazakhstan, grant AP14872061.
@@
N.U. and M.Z. conceived and designed the study. N.U. developed the methodology. M.Z. and S.A. implemented the software. M.Z. and A.A. conducted validation. S.A. performed formal analysis and data curation. N.U. led the investigation. S.A. provided resources. S.T. and M.Z. prepared the original manuscript draft. A.S. contributed to manuscript refinement and provided feedback on the writing. D.T. contributed to manuscript review and editing. M.Z. created the visualizations. N.U. supervised the project. All authors have read and approved the final manuscript

### v3

# The Role of Fractal Dimension in Wireless Mesh Network Performance

@@ Abstract
Wireless mesh networks (WMNs) depend on the spatial distribution of nodes, which directly influences connectivity, routing efficiency, and overall network performance. Conventional models typically assume uniform or random node placement, which inadequately represent the complex, hierarchical spatial patterns observed in practical deployments. In this study, we present a novel algorithm that constructs WMN topologies with tunable fractal dimensions, allowing precise control over spatial self-similarity. By systematically varying the fractal dimension, the algorithm generates network layouts spanning a continuum of spatial complexities, ranging from sparse fragmented clusters to dense, cohesive structures. Through NS-3 simulations, Key performance metrics including throughput, latency, jitter, and packet delivery ratio were evaluated across a range of fractal dimensions. Comparative evaluations against classical random, small-world, scale-free, grid and hierarchical tree networks models reveal that high-dimensional fractal topologies achieve enhanced resilience and throughput under equivalent conditions. These findings demonstrate the potential of fractal geometry as a design paradigm for scalable and efficient WMN architectures.
@@ Results
Figure 1 illustrates that at a low fractal dimension ($D=1$), the network fragments into sparsely connected components, forming isolated clusters with weak interconnections. This fragmentation results from the inherent spatial sparsity caused by the recursive process of low-dimensional generation. As the fractal dimension increases, the network becomes increasingly integrated and spatially uniform. At $D=2$, the topology resembles a grid-like tiling, while at $D=3$, more cross-linked central paths emerge, forming distinctive cruciform structures. Interestingly, from $D\in[4,8]$, the generated networks converge toward a stable topological regime characterized by dense, isotropic interconnectivity and robust multi-hop pathways. This suggests a structural saturation effect, where further increases in fractal dimension yield diminishing changes in spatial arrangement. When $D=7$ and continuing through $D=9$, ring-shaped or circular substructures begin to appear within the network topology. These formations represent the emergence of locally clustered subnetworks embedded within the overall topology, rather than a uniformly dispersed node distribution. This shift suggests that higher fractal dimensions promote localized aggregation, likely because of recursive spatial overlap during the point-generation process.
@@
Figure 3: Global efficiency, algebraic connectivity, and average edge-disjoint paths (EDP) versus fractal dimension $D=1$ to $D=10$. Unweighted networks show higher efficiency but lower connectivity, and EDP is shown for unweighted networks only
@@
Figure 5: Comparative performance analysis of fractal and baseline networks. Quality-of-service (QoS) metrics—including throughput, packet delivery ratio (PDR), end-to-end delay, and jitter are evaluated across fractal, scale-free (BA), small-world (WS), random (ER), grid and hierarchical tree topologies under the same simulation conditions.
To reveal the structural origins of these performance trends, topological indicators including global efficiency, algebraic connectivity, and average edge-disjoint paths were further examined (Figure 3). Global efficiency is defined as the average of the inverse shortest path lengths between all node pairs. At low fractal dimensions, $D=1$, the network is highly fragmented with low efficiency. For $D=2$ and $D=3$ , connectivity improves and the clusters become more spatially distributed, leading to a clear increase in efficiency, although the topology is still not fully stabilized. Around $D=4.5$, a topological transition point is reached: the network becomes spatially isotropic, and subsequent increases in $D$ lead to structurally similar but denser layouts, the network approaches a mature, cohesive topology, and efficiency plateaus, reflecting structural saturation. Interestingly, comparing weighted and unweighted networks reveals that the unweighted efficiency is consistently higher with a similar trend. This occurs because the weighted network penalizes long-range or weak links with lower weights, effectively increasing the “cost” of certain paths. In contrast, unweighted networks treat all links equally, allowing more alternative paths to contribute fully to efficiency, thereby producing higher average values.
Figure 6: Per-node load, coefficient of variation (CV), and Jain’s fairness index versus fractal dimension $D$. Mean load decreases with $D$, while CV and Jain’s index capture non-monotonic shifts in load balancing.
Algebraic connectivity, representing spectral robustness and derived from the Laplacian matrix weighted by link capacity, increases steadily between $D=1$ and $D=4.5$, reflecting improved interconnectivity and reduced path fragmentation. Beyond $D=4.5$, the growth plateaus, indicating structural saturation where further increases in $D$ yield marginal connectivity gains. A similar trend is observed for the unweighted network, though overall connectivity is lower because unweighted links ignore capacity differences, reducing the effective contribution of weak or long-range connections. This transition coincides with the stabilization of global efficiency and confirms that early topological shifts are the primary drivers of performance enhancement. To further quantify path diversity, we calculated the average number of edge-disjoint paths between node pairs in the unweighted network. EDPs grow up to $D=4.5$, reflecting increasing redundancy and alternative routing options, then plateau with small fluctuations due to the discrete nature of path counting. By Menger’s theorem, the EDPs between nodes is bounded by the minimum cut (or degree) [42], aligning with the observed saturation.
To evaluate load balancing beyond per-node averages, we first examined the total received bytes per node. While the per-node load alone shows a general trend of decreasing variance with increasing fractal dimension, a more detailed view using the coefficient of variation (CV) and Jain’s fairness index reveals complex fluctuations across $D$. As shown in Figure 6, the CV rises from $D=1$ to $D=2$, slightly decreases, increases up to $D=4.5$, then drops toward $D=7$, rises to $D=8$, briefly decreases, and increases again toward $D=9$. In contrast, Jain’s index follows an opposite but similarly non-monotonic pattern. Notably, in the intermediate range D=5.5D=5.5-$7.5$, the node load becomes relatively uniform and path diversity is high, resulting in peak network performance consistent with the throughput, latency, jitter, and packet delivery metrics as Figure 4. These observations highlight that while per-node averages provide a baseline, the combination of load distribution and path diversity drives optimal performance in specific fractal dimension ranges.
@@
To further contextualize the advantages of high-dimensional fractal topologies, networks were compared with several classical baselines including random geometric graphs, scale-free, small-world, grid and hierarchical tree configurations under equivalent traffic conditions (Figure 5). For this comparison, three representative fractal networks with dimensions $D=2$, $D=6.5$, and $D=7.5$ were selected. When $D=2$, the quasi-regular structure resembles grid-like arrangement. $D=6.5$ and $D=7.5$ were selected as representative high-performance configurations, based on the performance peak observed across all QoS metrics. These dimensions correspond to structural saturation points, beyond which further increases in D yield diminishing returns in performance. Including these values allowed us to benchmark near-optimal fractal networks against classical topologies. The results confirm that fractal-based networks, particularly those at higher $D$, consistently perform better than baseline networks. In particular, the performance of $D=2$ is nearly identical to that of the grid topology, indicating that low-dimensional fractal layouts can emulate deterministic spatial structures. While BA networks exhibit degree-based hierarchy and topological self-similarity, they lack the spatial self-similarity of our fractal networks. BA topologies form centralized hubs, whereas fractal networks display recursive spatial clustering, enabling better path diversity and load distribution. As a result, fractal networks consistently outperform BA in key performance metrics. However, hierarchical structures like BA still offer competitive performance, though slightly inferior due to their hub-centric bottlenecks under load. These findings demonstrate that a higher fractal dimension yields a dual benefit: improved structural cohesion and enhanced performance.
@@ Discussion
From a theoretical standpoint, WMN performance is governed by fractal geometry, connectivity and routing efficiency. For a node set of fractal dimension $D$, the mass–radius law $N(s)\propto s^{D}$ implies that with fixed range $r$, the mean degree scales as $\langle k\rangle\!\propto\!r^{D}$ [43, 44, 45]. Larger $\langle k\rangle$ increases clustering and reach, creating denser communities and redundant links. By Menger’s theorem, the number of edge-disjoint paths between two nodes is determined by the minimum cut, so higher degree yields more alternative routes [46]. Spectral theory complements this view: for the normalized Laplacian, algebraic connectivity $\lambda_{2}$ obeys
$\tfrac{1}{2}\phi(G)^{2}\leq\lambda_{2}\leq 2\phi(G),$ (1)
linking conductance $\phi(G)$ to resilience and expansion [47]. Global efficiency $E$ (Equation (8)), the mean inverse shortest-path length, measures multi-hop effectiveness. As $D$ increases, networks evolve from fragmented to efficient topologies. In our simulations ($N=85$, $r=250$ m), gains plateau for $D\gtrsim 4$, due to a cut bound path diversity limited by minimum degree $\delta(G)$, and a geometric bound fixed $N$ and $r$ restricting diameter and $E$. These bounds explain the saturation of $\lambda_{2}$, $E$, path diversity and fairness, while denser layouts still improve throughput, delivery ratio and delay. Thus fractal dimension acts as a tuning parameter shaping connectivity, routing geometry and performance. The interplay between load distribution and path diversity explains why peak performance occurs around $D\approx 6.5$ rather than at higher dimensions. In this intermediate regime, connectivity and efficiency are already high, while traffic remains well-balanced across nodes, as reflected in low CV and high Jain’s index (Figure 6). At larger $D$ ($\geq 8$), new links primarily add redundant neighbors without creating independent routes: the cut bound limits and the geometric bound limits efficiency. As a result, flows concentrate on highly connected nodes, raising CV and lowering fairness, which constrains throughput and latency improvements despite higher density. This explains the alignment betweenFigures 3and 6, and shows that achieve near-optimal network performance emerges when structural connectivity and balanced load coincide.
@@
This section presents an algorithm that constructs WMN topologies by explicitly setting the fractal dimension $D$ as an input parameter to control the spatial distribution of the nodes. The nodes are positioned according to the specified fractal dimension, and the edges of the network are formed by connecting pairs of nodes within a fixed communication radius $r$ . The resulting fractal-based topologies are then analyzed using NS-3 simulations, where key performance indicators such as throughput, end-to-end delay, jitter, and PDR are measured under uniform traffic scenarios. To implement the proposed algorithm, the following steps are performed:
@@
[…]
@@
1:Input: Target number of nodes NN, fractal dimension $D$
2:Output: First NN node coordinates $(x_{i},y_{i})$
3:Initialize empty list: $\mathit{points}\leftarrow\emptyset$
4:Compute maximum recursion depth: $\mathit{depth}_{\max}\leftarrow\left\lfloor\frac{\log(3N)}{\log 4}\right\rfloor$
@@
16:Generate($x_{0},y_{0},s_{0},\mathit{depth}_{\max}$)
17:return First NN points from $\mathit{points}$
@@
To establish a rigorous baseline for evaluating the structural and communication advantages of fractal-based networks, three classical generative network models were constructed: Erdős–Rényi (ER), Watts–Strogatz (WS), and Barabási–Albert (BA), all generated under consistent conditions of network size and mean degree. Specifically, ER networks were generated using the $G(n,p)$ model with $n=85$ nodes and edge probability $p=\frac{k}{n-1}$, targeting a mean degree $k=4.5$. WS networks were initialized as regular ring lattices with each node connected to its $k=4$ nearest neighbors and then rewired with a probability of $p=0.1$ to introduce small-world properties. BA networks were created using a preferential attachment mechanism, starting with a small fully connected seed and sequentially adding nodes, each establishing $m=2$ links based on the degree of the existing node. To ensure a valid comparison with the fractal networks, all generated networks were required to be fully connected, with repeated sampling applied as needed. Random seeds were fixed to maintain reproducibility and consistent layout configurations. In addition to the classical ER, WS, and BA networks, we introduced two deterministic topologies to broaden the structural comparison: a regular grid network and a hierarchical tree network. The grid network was constructed as a $9\times 9$ two-dimensional lattice ($81$ nodes), augmented by $4$ additional nodes placed at the centers of the four corner cells, resulting in a total of $85$ nodes.The grid layout provides a spatially uniform and deterministic structure with regular local neighborhoods, emulating idealized deployments in structured environments such as factory floors or urban grids. The hierarchical network was generated as a balanced tree with branching factor $r=4$ and height $h=3$, yielding exactly $85$ nodes. Each non-leaf node connects to $4$ child nodes in a recursive structure, forming a layered topology that emphasizes centralized hierarchy and minimal path diversity.
These baseline models represent, respectively, random (ER), small-world (WS), scale-free (BA), spatially regular (grid), and hierarchically modular (tree) regimes commonly studied in the network topology literature. They serve as essential structural archetypes for evaluating the topological distinctiveness and impact on performance of high-dimensional fractal geometries. The corresponding topologies are illustrated in Figure 8. This diversity provides a set of contrastive regimes for benchmarking the behavior of fractal-based network designs.
Figure 8: Visualization of classical network models (ER, WS, BA,grid and hierarchical tree) used as baseline topologies for comparison. All configurations have identical node counts and mean degrees to ensure fair evaluation. Node colors represent average link capacity, highlighting local connectivity patterns.
@@
The simulation environment consisted of 85 static nodes deployed using the proposed spatially heterogeneous, self-similar fractal layout across a $500\,\mathrm{m}\times 500\,\mathrm{m}$ area. The node positions were generated using the proposed algorithm (Algorithm 1), with the fractal dimension $D\in[1,10]$. The resulting self-similar and spatial distributions were visually verified.
@@
In addition to the standard metrics, including throughput, end-to-end delay, jitter, and packet delivery ratio (PDR) obtained via NS-3’s FlowMonitor module, we evaluated additional topological and load-balancing indicators to quantify path diversity and traffic distribution.
The global efficiency of the network is defined as the average of the inverse shortest path lengths between all node pairs:
$E=\frac{1}{N(N-1)}\sum_{i\neq j}\frac{1}{d_{ij}},$ (8)
where $d_{ij}$ denotes the shortest path length between nodes $i$ and $j$, considering weighted or unweighted graphs as specified.
Algebraic connectivity $\lambda_{2}$ is computed as the second-smallest eigenvalue of the graph Laplacian matrix
L=DA,L=D-A, (9)
where $D$ is the degree matrix and $A$ is the adjacency matrix, weighted by link capacity for the weighted case.
The average number of edge-disjoint paths (EDPs) between node pairs is obtained using a max-flow algorithm (e.g., Ford-Fulkerson) on the unweighted graph and then averaged over all pairs.
To assessload balancing, we analyzed the total received bytes per node. The coefficient of variation (CV) for node loads $x=(x_{1},\dots,x_{N})$ is defined as
$\mathrm{CV}=\frac{\sigma}{\mu},$ (10)
where $\sigma$ and $\mu$ represent the standard deviation and mean of the node loads, respectively.
Jain’s fairness index $J$ is defined as [50]:
$J=\frac{\left(\sum_{i=1}^{N}x_{i}\right)^{2}}{N\sum_{i=1}^{N}x_{i}^{2}},$ (11)
quantifying the fairness of resource allocation, where $J=1$ corresponds to perfect fairness and $J=1/N$ indicates the worst case. Note that CV and $J$ are related by
$J=\frac{1}{1+\mathrm{CV}^{2}},$ (12)
explaining their inverse fluctuations.
Data availability
The simulation data generated and analysed during this study are available at the following GitHub repository: https://github.com/Zaidyn-marat/Fractal-based-WMNs-Topology. The repository includes the full fractal topology generator, NS-3 simulation scripts, and performance evaluation code used to produce the results presented in this paper. Additional datasets are available from the corresponding author upon reasonable request.
@@ Acknowledgements
We would like to express our sincerest gratitude to the Al-Farabi Kazakh National University for supporting this work by providing computing resources (Department of Physics and Technology).
Funding
This research was funded by the Committee of Science of the Ministry of Science and Higher Education of the Republic of Kazakhstan (Grant No. AP19674715).
@@
N.U. and M.Z. conceived and designed the study. N.U. developed the methodology. M.Z. and S.A. implemented the software. M.Z. and A.A. conducted validation. S.A. performed formal analysis and data curation. N.U. led the investigation. S.A. provided resources. S.T. and M.Z. prepared the original manuscript draft. A.S. contributed to manuscript refinement and provided feedback on the writing. D.T. contributed to manuscript review and editing. M.Z. created the visualizations. N.U. supervised the project. All authors have read and approved the final manuscript.